# Enhanced sensing of Stark weak field under the influence of Aubry-André-Harper criticality

### v2

@@ VI Conclusions
In the realm of quantum many-body systems, achieving super-Heisenberg limit precision remains a formidable challenge. This paper focuses on engineering efficient quantum many-body sensors. In an excellent recent new work it has been proposed that efficient weak field sensors can be designed by utilizing Stark localization as a resource [43]. Specifically, it has been demonstrated that under Stark localization, the quantum Fisher information (QFI) scales as $L^{6}$, i.e., a beyond-Heisenberg scaling is achieved.
@@
This work proposes to use different classes of localization inducing potentials that can enhance the precision of parameter estimation even further. Near the criticality of the AAH potential, better precision in estimating weak fields of Stark strength can be achieved with the QFI scaling as $L^{6.7}$ in comparison to the pure Stark system for which the QFI is reported to scale as $L^{6}$. In addition, this work proposes an experimentally accessible observable, the charge-density-wave (CDW) order denoted as $\hat{O}_{cdw}$. Remarkably, this observable also exhibits super-Heisenberg scaling with the QFI scaling as $F_{cdw}\sim L^{6.2}$. Moreover, another observable, $\hat{O}_{H_{2}}$, has been identified for which the associated OFI not only saturates to the QFI but also matches the scaling behavior of the QFI. For many-body systems, particularly under the condition of half-filling, the scaling of the QFI with system size is found as $F_{Q}\sim L^{6.6}$. Similarly, for the half-filled case, the OFI corresponding to the operator $\hat{O}_{H_{2}}$ exhibits the same scaling behavior as the QFI, that is $F_{H_{2}}\sim L^{6.6}$.
@@
In summary, our work proposes an experimentally realizable protocol for achieving enhanced sensing of the Stark weak field by invoking another type of localization-inducing potential in the form of the AAH potential and by exploiting the AAH criticality. One may, in principle, envision the underlying strength of the concept which can, in principle, be generalized for criticality-based quantum sensors – In the context of precision estimation of an unknown parameter within the QMB platform, it might be advantageous to introduce additional control parameters, that, in turn, introduces new kinds of quantum criticality, for obtaining enhanced quantum advantage by utilizing the competing effects of multiple critical points.

### v3

@@ VIII Conclusions
In the realm of quantum many-body systems, achieving super-Heisenberg limit precision remains a formidable challenge. This paper focuses on engineering efficient quantum many-body sensors. In an excellent recent new work it has been proposed that efficient weak-field sensors can be designed by utilizing Stark localization as a resource [46]. Specifically, it has been demonstrated that under Stark localization for ground state probe, the quantum Fisher information (QFI) scales as $L^{6}$ and for midspectrum state the QFI scales as $L^{4.1}$, i.e., a beyond-Heisenberg-limit scaling is achieved.
@@
This work proposes to use different classes of localization-inducing potentials that can enhance the precision of parameter estimation even further. Near the criticality of the AAH potential, better precision in estimating weak fields of Stark strength can be achieved with the QFI scaling as $L^{6.7}$ for ground state, and $L^{5.6}$ for the midspectrum state, respectively. In comparison, for the pure Stark system, the QFI is reported to scale as $L^{6}$ and $L^{4.1}$ for ground state and midspectrum state, respectively. In addition, this work examines the charge-density-wave (CDW) order, $\hat{O}_{cdw}$, which is an experimentally accessible observable. This observable also exhibits super-Heisenberg scaling with the OFI scaling as $F_{cdw}\sim L^{6.2}$ for ground state as a probe. Moreover, another observable, $\hat{O}_{H_{2}}$, has been identified for which the associated OFI not only saturates to the QFI but also matches the scaling behavior of the QFI. For the midspectrum state, the OFI for CDW scales with $L$ as $F_{cdw}\sim L^{5}$, and the scaling exponent of OFI for $\hat{O}_{H_{2}}$ saturates the scaling of the QFI. Now, for the thermal state as a probe, we have shown that the QFI saturates to the HL for all ranges of the temperature considered in this work when the system is prepared in the delocalized phase ($h=10^{-8}$). When the system is at localized phase $h=0.05$, the QFI collapses for all system sizes at the low temperature limit, i.e., the system remains scale invariant and there is no scaling with system sizes. Still, at higher temperatures the QFI saturates to the HL. However, as expected, the absolute values of the QFI and OFI in the high-temperature regime are reduced compared to those at low temperatures, indicating that the probes are more effective when thermal fluctuations are minimal. Further studies have been performed for the system at half-filling, for which the scaling of the QFI with system size for the ground state is found to be $F_{Q}\sim L^{6.6}$. Similarly, for the half-filled case, the OFI corresponding to the operator $\hat{O}_{H_{2}}$ exhibits the same scaling behavior as the QFI, that is $F_{H_{2}}\sim L^{6.6}$.
@@
In summary, our work proposes an experimentally realizable protocol for achieving enhanced sensing of the weak Stark field by invoking another type of localization-inducing potential in the form of the AAH potential and by exploiting the AAH criticality. One may, in principle, envision the underlying strength of the concept which can, in principle, be generalized for criticality-based quantum sensors – In the context of precision estimation of an unknown parameter within the QMB platform, it might be advantageous to introduce additional control parameters, that, in turn, introduces new kinds of quantum criticality, for obtaining enhanced quantum advantage by utilizing the competing effects of multiple critical points.